OpenStudy (anonymous):

Show an equation and a solution for the problem. Crane A can unload a ship in 10 hours and crane B can unload it in 14 hours. How long will it take the two cranes to unload the ship working together? Round to the nearest tenth.

OpenStudy (paxpolaris):

Crane A can unload a ship in 10 hours so, A uploads 1/10 of the ship in 1hr

OpenStudy (paxpolaris):

B unloads 1/14 of the ship in 1hr

OpenStudy (paxpolaris):

if x is the number of hrs it takes working together. \[\Large x \left( \frac 1 {10} +\frac 1 {14} \right)=1\]

OpenStudy (anonymous):

thanks!!!!!